OpenStudy (anonymous):

the length of a rectangle is 12 ft longer than twice the width. If the perimeter is 90 ft, find the length and the width of the rectangle.

OpenStudy (anonymous):

2l+2w=90

OpenStudy (anonymous):

2w*12=90 ?

OpenStudy (alexwee123):

solve out for x

OpenStudy (campbell_st):

draw a picture then 2(2w + 12) + 2w = 90 solve the equation to find the width